OpenStudy (anonymous):

If f(x) = x^ 2 + 1 and g(x) = 3x + 1, find 2f(1) + 3g(4).

OpenStudy (anonymous):

f(1)= 1^2 + 1 = 2 2f(1)= 4 g(4)= 3*4 + 1= 13 3g(4)= 39 Ans: 4 + 39 = 43

OpenStudy (anonymous):

thanks